What volume is occupied by 0.869 mol of CO2 at 214.2 K and 961 mmHg?
Answer #1

we have:

P = 961 mm Hg

= (961/760) atm

= 1.2645 atm

n = 0.869 mol

T = 214.2 K

we have below equation to be used:

P * V = n*R*T

1.2645 atm * V = 0.869 mol* 0.0821 atm.L/mol.K * 214.2 K

V = 12.1 L

Answer: 12.1 L